### v3

# Universal relation between $C_{T}$ and the CFT Weyl anomaly

Abstract
We establish a universal relation between the coefficient $C_{T}$ of the energy–momentum tensor two‑point function and the coefficient $c$ multiplying the term quadratic in the Weyl tensor in the Weyl anomaly of a generic even‑dimensional conformal field theory. Our first derivation combines long‑known holographic results for $C_{T}$ and for the Weyl anomaly in Einstein bulk gravity with a recently obtained Chern–Gauss–Bonnet formula for compact Einstein manifolds. This theorem isolates the Weyl‑squared contribution in the relation between the Euler density and the $Q$-curvature, allowing us to identify the relevant quadratic term unambiguously. We then provide two genuine CFT derivations based on the renormalization‑group running of the TT correlator with respect to the arbitrary but necessary mass scale $\mu$. Several known examples are revisited to illustrate and validate the general result.
pacs:
04.50.+h, 04.70.Bw
I Introduction
Conformal field theories (CFTs) occupy a central position in modern theoretical physics. They describe quantum field theories that are invariant under the full conformal group, which extends Poincaré symmetry by including scale transformations and special conformal transformations. This enlarged symmetry strongly constrains the structure of correlation functions, the operator spectrum, and the renormalization‑group behavior of the theory. In any dimension $d$, the local operators of a CFT organize into irreducible representations of the conformal group, labeled by their scaling dimensions and Lorentz spins. Among these, the energy–momentum tensor $T_{ab}$ plays a distinguished role: it generates spacetime symmetries and its correlation functions encode universal dynamical data. In particular, the two‑point function of $T_{ab}$ is fixed up to an overall coefficient $C_{T}$, which serves as a measure of the number of degrees of freedom of the theory and appears in a variety of physical contexts, from entanglement entropy to conformal collider bounds. In our conventions, the two-point function of the energy-momentum tensor in flat space reads
$\langle T_{ab}(x)\,T_{cd}(0)\rangle=\frac{C_{T}}{S_{d}^{2}}\cdot\frac{\mathcal{I}_{ab,cd}(x)}{|x|^{2d}},$ (1)
where $S_{d}=2\pi^{d/2}/\Gamma(d/2)$ is the area of the unit (d1)(d-1)-sphere, ${\cal{I}}_{ab,cd}(x)$ represents the inversion operator on symmetric traceless tensors $\mathcal{I}_{ab,cd}=(I_{ac}I_{bd}+I_{ad}I_{bc})/2-\delta_{ab}\delta_{cd}/d$, which is made of the inversion tensor $I_{ab}=\delta_{ab}-x_{a}x_{b}/|x|^{2}$, and $|x|$ denotes the Euclidean distance between the two points (see, e.g., [1]).
In even spacetime dimensions, CFTs exhibit a trace (or Weyl or conformal) anomaly when coupled to a background metric [2]. Although the classical theory is Weyl invariant, the quantum effective action acquires a non‑vanishing trace proportional to local curvature invariants. The anomaly takes the schematic form [3, 4, 5]
$(4\pi)^{d/2}\langle T\rangle=-a\,E_{d}+\sum_{i}c_{i}I_{i}+ttd,$ (2)
where the Euler density $E_{d}$ carries the type-A central charge $a$, the pointwise Weyl invariants $I_{i}$ carry the type-B central charges $c_{i}$, and ttdttd stands for trivial anomalies or trivial total derivatives whose conformal primitives are local curvature invariants. The central charges, traditionally denoted $a$ and $c$ in four dimensions and by suitable generalizations in higher even dimensions, are intrinsic CFT data.
The trace anomaly anchors RG irreversibility, fixes stress‑tensor data in CFTs, generates (non-local) quantum effective actions, controls vacuum/Casimir/Hawking effects in curved space, and shapes quantum corrections in cosmology and black‑hole physics, among many other physical applications in quantum gravity, black hole physics, inflationary cosmology, string theory and statistical mechanics (see e.g. [6] for a review).
The connection between $C_{T}$ and the Weyl anomaly coefficients exhibits a rich and intricate structure. Although these quantities arise from different physical considerations – flat‐space correlation functions versus curved‑space anomalies – they are not independent. In holographic theories, for example, both can be computed from the same bulk gravitational action, leading to universal relations. More generally, following the lead of Osborn and Petkou [1], since the integrated trace anomaly determines the running of the CFT partition function with the mass scale $\mu$, then this ties the Weyl‑squared term in the anomaly to the normalization of the stress‑tensor two‑point function around flat space. So that in any even $d$, $C_{T}$ is fixed by the part of the Weyl anomaly that is quadratic in the metric fluctuation, everything else—higher powers of the Weyl tensor, trivial total derivatives, topological terms—does not feed into $C_{T}$. Here is how that plays out dimension by dimension:
  • d=2: There is only one central charge, $c$, and it is defined by
    $4\pi\langle T\rangle=\frac{c}{6}R.$
    This controls both the trace anomaly and the 2‑point function given by
    $C_{T}=2\,c\quad(d=2).$
    This result, of course, complies with the well-known results for the conformal scalar $C_{T}=\frac{d}{d-1}=2$ and the convention that for the free boson $c=1$.
  • d=4: The anomaly is determined by $a$ and $c$, coefficients of the Euler density and Weyl tensor squared terms, respectively, modulo a trivial total derivative (ttd)
    $(4\pi)^{2}\langle T\rangle=-a\,E_{4}\,+\,c\,W^{2}\,+\,ttd\,.$
    Here, $C_{T}$ is uniquely proportional to the $c$ coefficient associated with the square of the Weyl tensor. The 2‑point function of $T_{ab}$ is controlled by the coefficient of $W^{2}$, i.e., by $c$. The Euler term $E_{4}$ and the trivial total derivative do not contribute at quadratic order in the metric fluctuation around flat space, and, as shown in [1],
    $C_{T}=160\,c\quad(d=4)\,.$
  • d=6: As $d$ increases, the number of independent Weyl invariants ($I_{i}$) grows. In $d=6$, there are three such coefficients ($c_{1},c_{2},c_{3}$). Of the three Weyl invariants, only one relates to $C_{T}$. The anomaly is usually written as
    $(4\pi)^{3}\langle T\rangle=-a\,E_{6}+c_{1}I_{1}+c_{2}I_{2}+c_{3}I_{3}+ttd,$
    where $I_{1}$ and $I_{2}$ are cubic in the Weyl tensor, while $I_{3}$ contains terms of the schematic form $W\nabla^{2}W+lot+ttd$. The lower-order terms in derivatives (lotlot) clearly do not contribute to the quadratic expansion around flat space and are not directly related to $C_{T}$. Therefore, it is the central charge $c_{3}$ that relates to $C_{T}$ [1, 7]
    $C_{T}=3024\,c_{3}\quad(d=6)\,.$
  • d=8: The pattern expands even further. There is a larger number of Weyl invariants $I_{i}$ in the anomaly. Fortunately, only those invariants whose expansion contains quadratic terms of the Weyl tensor (with derivatives allowed) can contribute to $C_{T}$. Strictly cubic and higher curvature invariants start only at cubic order in the metric fluctuation and thus only affect 3‑point and higher stress‑tensor correlators.
    In holographic Einstein gravity, where the bulk action is just the Hilbert-Einstein action plus a negative cosmological constant, the coefficients of the holographic Weyl anomaly [8, 9] are all determined by a single bulk coupling: that of the critical $Q$-curvature [10]. That enforces linear relations among the various $c_{i}$ and $a$, but structurally, $C_{T}$ still tracks the unique quadratic-in-$W$ direction in the space of invariants, just as in 6d it tracks $c_{3}$. In $8d$, Chen and Lü [11] (see also [12, 13]) have recently found a quadratic in Weyl invariant $I^{(8)}_{10}$ whose coefficient $c_{10}$ becomes equal to $a$ in holographic Einstein gravity. Comparing with the long-known result of Liu and Tseytlin for $C_{T}$ [14], they obtain (in our conventions)
    $C_{T}=69120\,c_{{}_{10}}\quad(d=8)\,.$
  • d>>2, even: In higher dimensions, the precise structure of the type-B Weyl invariants is largely unknown, but for our present purposes, we can safely focus on the term quadratic in the Weyl tensor. So that, up to higher curvature terms and trivial total derivatives, we have
    $(4\pi)^{d/2}\langle T\rangle=-a\,E_{d}+c\,\frac{d}{4}W^{(d)}_{2,1}+lot+ttd,$ (3)
    where we have chosen a particular pointwise Weyl invariant $W^{(d)}_{2,1}$, recently introduced by Case et at. in [13] that descends from the ambient $\left(\widetilde{\nabla}^{2}\right)^{d/2-2}\widetilde{W}^{2}$ in the Fefferman–Graham program. Restricted to Einstein manifolds, it enters an explicit expression for the Chern–Gauss–Bonnet formula that allows for the interchange between the Euler density and the $Q$-curvature. With this insight, we tune the coefficient $c$ in the anomaly to enforce the equality $c=a$ when the anomaly becomes just the critical $Q$-curvature.
    Alternatively, we are also free to choose instead a different Weyl invariant quadratic in the Weyl tensor that naturally extends the $W^{2}$ and the $I_{3}$ from 4d and 6d, respectively (see e.g. [15]). The numerical coefficient is obtained by the expansion of the $Q$-curvature to terms quadratic in curvature (see subsection II.C below) to enforce the equality $c=a$ for the $Q$-curvature,
    $(4\pi)^{d/2}\langle T\rangle=-a\,E_{d}+c\,\frac{d(d-2)}{8(d-3)}\left[W\left(\nabla^{2}\right)^{d/2-2}W+lot\right]+ttd\quad.$ (4)
In what follows we will explain how to obtain the following universal relation between $C_{T}$ and $c$, that accommodates the low dimensions and extrapolates to any even higher dimension [^1] Dimension two is special, our general result can be modified to include this particular case as well by absorbing a factor six in the definition of the 2d $c$, setting instead $4\pi\langle T\rangle=c\,R$
$C_{T}=\dfrac{d}{d-1}\cdot\dfrac{(d+1)!}{(d/1)!}\cdot c\qquad\text{ for even }d>2.$ (5)
Alternatively, in terms of the critical $Q$-curvature $Q_{d}$ and the coefficient $C_{TT}$ of the TT-correlator in momentum space  [^2] We thank P. McFadden for pointing out the suitability of the $C_{TT}$ coefficient, as opposed to the more conventional $C_{T}$., the explicit dimension dependence can be absorbed and the universality appears in a surprisingly simple form
$\boxed{\langle T\rangle\,=\,2\,C_{TT}\cdot Q_{d}+lot+ttd,}$ (6)
where the lower-order terms in derivatives (lotlot) now include the Euler density and only the pointwise Weyl invariants of the type-B trace anomaly which are cubic and higher in curvature.
Understanding these relations is essential for characterizing the space of consistent CFTs in higher dimensions and to connect flat‑space observables with geometric responses. The present work develops this connection in full generality for even‐dimensional CFTs. We follow a holographic approach and combine the long-known holographic results for $C_{T}$ [14] and Weyl anomaly [8, 9, 10] for bulk Einstein gravity, together with the recently derived Chern-Gauss-Bonnet theorem on a compact Einstein manifold that allows one to pinpoint the term quadratic in the Weyl tensor in the relation between Euler density and $Q$-curvature. We also confirm our findings through two purely CFT computations that exploit the fact that the trace anomaly determines the scaling behavior of the CFT partition function and effective action.
We start in Section II.1, by reviewing the holographic computation of $C_{T}$ and relating it to the bulk Newton’s constant, or, equivalently, the Planck length. In II.2, we discuss the holographic Weyl anomaly for Einstein gravity with a negative cosmological constant and its relation with the $Q$-curvature. In II.3, we present the Chern-Gauss-Bonnet theorem on compact even-dimensional Einstein manifolds, as recently obtained by Case et al. [13], which relates the Euler density to the $Q$-curvature and pointwise Weyl invariants. In particular, we keep track of the coefficient $c$ of the quadratic term in the Weyl tensor. In II.4, we combine the previous ingredients to obtain the universal relation between $C_{T}$ and $c$. In Section III, we present two alternative derivations within the CFT. Finally, in Section IV, we compare the universal relation with several known examples, illustrate them, and conclude in Section V.
II Universal relation between $C_{T}$ and $c$ from holographic data
We first determine a universal relation by examining holographic data obtained from two distinct computations. On the one hand, the holographic computation of $C_{T}$ in bulk Einstein gravity involves the AdS background radius and the Planck length. On the other hand, we report the coefficient of the quadratic term in the Weyl tensor of the corresponding holographic Weyl anomaly, which is essentially given by the $Q$-curvature. We then make use of the Chern-Gauss-Bonnet theorem of Case et al. [13] on compact Einstein manifolds to pinpoint the precise relation between $Q$-curvature and the Weyl invariant relevant for our analysis. This coefficient is expressed in terms of the AdS radius and Planck’s length. Eliminating the two-dimensional quantities, we end up with the algebraic relation between the two coefficients of the CFT data.
II.1 Holographic derivation of $C_{T}$ from bulk Einstein gravity
One of the early successes of the AdS/CFT calculational prescription was the holographic derivation of $C_{T}$ for the CFT dual to bulk Einstein gravity by Liu and Tseytlin [14]. The (Euclidean) gravitational action is given by
[…]
where $l_{\text{P}}$ is the Planck length, the bulk cosmological constant can be written in terms of the radius of the vacuum AdS solution $\hat{\Lambda}=-d(d-1)/L^{2}$, the boundary terms comprise the standard Gibbons–Hawking–York term [18] and an effective boundary cosmological constant $\hat{\lambda}$ that is tuned later on to (d1)/L(d-1)/L to ensure the conformal invariance of the action computed on the solution of the Dirichlet problem for the metric fluctuation. The quadratic part of the gravitational action, as a function of the boundary metric fluctuation $h$, turned out to be given by a nonlocal kernel that matches the CFT expectation for the two-point function of the energy-momentum tensor in flat space
$S^{(2)}_{\text{EH}}[h]=\dfrac{\Gamma(d+2)\,L^{d-1}}{8\pi^{d/2}(d-1)l_{\text{P}}^{d-1}}\int_{\mathbb{R}^{d}}d^{d}x\int_{\mathbb{R}^{d}}d^{d}y\dfrac{h_{ab}(x)\,\mathcal{I}_{ab,cd}(x-y)\,h_{cd}(y)}{|x-y|^{2d}}~.$ (8)
Performing the second functional variation with respect to the source $h$, one obtains the holographic two-point correlation function for the energy-momentum tensor and, with it, the $C_{T}$ coefficient
$C_{T}=4\pi^{d/2}\cdot\frac{\Gamma(d+2)}{(d-1)\,\Gamma^{3}(d/2)}\cdot\frac{L^{d-1}}{l_{\text{P}}^{d-1}}.$ (9)
II.2 Holographic Weyl anomaly for bulk Einstein gravity and $Q$-curvature
As an outgrowth of the IR/UV connection[19], bulk infrared (IR) divergences can be mapped to the ultraviolet (UV) divergences of the boundary theory, and thus, the Weyl anomaly of the dual CFT are determined from the bulk action. The detailed computation of this holographic Weyl anomaly was originally carried out by Henningson and Skenderis [8, 9]. This sparked interest among conformal geometers and led Graham and Zworski [10] to discover that the integrated holographic Weyl coincides with the integrated $Q$-curvature – a central object in Conformal Geometry.
There is a particular bulk metric, a solution to the Einstein equation, that enables straightforward computations and allows us to identify the numerical coefficients in front of the $Q$-curvature. Consider the particular Poincaré-Einstein metric $\hat{g}$ with an Einstein $g_{E}$ metric at its conformal infinity[20]
$\hat{g}=L^{2}\dfrac{dx^{2}+(1-\lambda\,x^{2}/2)^{2}g_{{}_{E}}}{x^{2}}.$ (10)
Here $\lambda=R/2d(d-1)$ is a multiple of the necessarily constant Ricci scalar of the boundary Einstein manifold. The volume anomaly is given by the logarithmically divergent term of the regularized volume expansion
$\mbox{Vol}[\{x>\epsilon\}]=L^{d+1}\,vol[g_{{}_{E}}]\int_{\epsilon}\dfrac{dx}{x^{d+1}}(1-\lambda\,x^{2}/2)^{d}\sim L^{d+1}\,vol[g_{{}_{E}}]\int_{\epsilon}\frac{dx}{x}(-\lambda/2)^{d/2}\binom{d}{d/2}$. (11)
The volume anomaly times the Einstein-Hilbert action on the PE/E metric, $d/L^{2}l_{{P}}^{d-1}$, gives the integrated holographic Weyl anomaly, so that one can read
$(4\pi)^{d/2}\langle T\rangle=-\dfrac{4(-\pi)^{d/2}}{\Gamma^{2}(d/2)}\,\dfrac{L^{d-1}}{l_{{P}}^{d-1}}\cdot Q_{d}+ttd,$ (12)
where we have used the fact that the critical $Q$-curvature, in our conventions  [^3] In particular, on the standard round unit sphere we have $Q_{d}=(d-1)!$., on a compact Einstein manifold reduces to $Q_{d}=(2\lambda)^{d/2}\cdot(d-1)!$. We need to establish the connection between the critical $Q$-curvature and the pointwise Weyl-invariant quadratic in the Weyl tensor.
II.3 Weyl-squared content of the $Q$-curvature
A key result in [13] is that any scalar contraction of the Weyl tensor on an Einstein manifold can be completed with total derivative terms, powers of the Laplacian, thereby making it a pointwise conformal invariant.
In particular, when applied to the Pfaffian that is polynomial in the Weyl tensor, one ends up with an explicit formula for the Pfaffian in terms of pointwise conformal invariants and a pure curvature scalar power that corresponds to the critical $Q$-curvature.
For our purposes, it is enough to focus on the connection between $E_{d},\,Q_{d}$ and $W^{(d)}_{2,1}$. We consider the following excerpt of Theorem (1.2) in [13]
$\mbox{Pf}=(2\lambda)^{d/2}\cdot(d-1)!!\,+\,\dfrac{(-2)^{2-d/2}}{(d/2-1)!}\cdot\frac{1}{8}\,W^{(d)}_{2,1}+lot+ttd.$ (13)
The convention for the Pfaffian is that the volume integral on a compact $d$-manifold produces $(2\pi)^{d/2}$ times the Euler characteristics. Therefore, to translate to our convention for the Euler density we first note that
$E_{d}=\dfrac{(-)^{d/2}}{2^{d/2}}\,\epsilon_{d}\,\epsilon_{d}\,Riem...Riem\qquad\text{ and }\qquad\mbox{Pf}=\dfrac{(d-1)!!}{2^{d/2}\,d!}\,\epsilon_{d}\,\epsilon_{d}\,Riem...Riem\,=\,\dfrac{(-)^{d/2}\,(d-1)!!}{d!}\cdot E_{d}~.$ (14)
Second, we recall that the critical $Q$-curvature on the compact Einstein is simply given by a multiple of the appropriate power of the constant scalar curvature
$Q_{d}=(2\lambda)^{d/2}\cdot(d-1)!.$ (15)
Thus, the $Q$-curvature is related with our preferred pointwise Weyl invariant  [^4] Note that our convention slightly differs from that of Case et al. [13]. We reserve $W^{(d)}_{2,1}$ to denote the $d$-dimensional invariant induced by the corresponding ambient curvature invariant. quadratic in the Weyl tensor $W^{(d)}_{2,1}$ as
$Q_{d}=\dfrac{(-1)^{d/2}}{d}\,\left\{E_{d}\,-\,\dfrac{d}{4}\,W^{(d)}_{2,1}\,+\,lot\,+\,ttd\right\}~.$ (16)
We are now in a position to write down the holographic value for $c$ from eqn.12
$c\,=\,a\,=\,\dfrac{d\,\pi^{d/2}}{[(d/2)!]^{2}}\,\dfrac{L^{d-1}}{l_{P}^{d-1}}.$ (17)
Alternatively, we can also connect the $Q$-curvature with a term quadratic in the Weyl tensor by careful examination of its leading and subleading terms in curvature (see discussion around eqn.(40))
$Q_{d}=-\,\dfrac{d-2}{8(d-3)}\,W(-\nabla^{2})^{d/2-2}W+lot+ttd.$ (18)
II.4 $C_{T}$ vs. $c$ from holographic data
We can now compare the two holographic results for $C_{T}$ and $c$ and eliminate the dependence on the AdS radius and Planck length to obtain
$\boxed{C_{T}=\frac{d}{d-1}\cdot\frac{(d+1)!}{\left(d/2-1\right)!}\cdot c}$ (19)
When written in terms of the $C_{TT}$ coefficient of the TT-correlator in momentum space (cf. [23])
$\langle\langle T_{ab}({\bf p})\,T_{cd}(-{\bf p})\rangle\rangle\,=\,C_{TT}\,\Pi_{ab,cd}({\bf p})\,p^{d},$ (20)
with $\Pi_{ab,cd}$ being the transverse-traceless projector and the Dirac delta enforcing momentum conservation factored out, the above relation simplifies to
$\boxed{(4\pi)^{d/2}\,C_{TT}=-\frac{(-1)^{d/2}\;d}{2}\cdot c}~.$ (21)
Without a genuine CFT derivation, the relation above might appear to be nothing more than a numerical coincidence. We therefore turn to the task of establishing this central relation by examining more closely the connection between the two‑point function of the energy–momentum tensor and the trace anomaly in a generic even‑dimensional CFT.
III Universal relation between $C_{T}$ and $c$ : CFT derivations
The CFT derivations exploit the fact that the trace anomaly controls the scale variation of the partition function $Z_{CFT}$ or effective action $W_{CFT}$ (see e.g. [1])
$\mu\dfrac{d}{d\mu}W_{CFT}=\int_{{\cal M}_{d}}dvol_{g}\,\langle T\rangle~.$ (22)
Upon functional differentiation with respect to the background metric, it is obtained the running of the energy-momentum tensor two-point function on flat space for the LHS of the above relation
$\mu\dfrac{d}{d\mu}\langle T_{ab}(x)\,T_{cd}(0)\rangle=\dfrac{C_{T}}{S_{d}^{2}}\,\mu\dfrac{d}{d\mu}\,\frac{{\cal{I}}_{ab,cd}(x)}{|x|^{2d}}=\dfrac{C_{T}}{S_{d}^{2}}\,\dfrac{\Delta^{T}_{ab,cd}}{4(d-2)^{2}d(d+1)}\,\mu\dfrac{d}{d\mu}\,\frac{1}{|x|^{2d-4}}~.$ (23)
Here $\Delta^{T}_{ab,cd}$ stands for a quartic differential operator that ensures conservation and tracelessness: $\Delta^{T}_{ab,cd}=(S_{ac}S_{bd}+S_{ad}S_{bc})/2-S_{ab}S_{bc}/(d-1)$, with $S_{ab}=\partial_{a}\partial_{b}-\delta_{ab}\partial^{2}$.
The scale dependence of $1/|x|^{2d-4}$ comes from the regularization prescription. With the following two identities within differential regularization [24]
$(i)$ $\quad\dfrac{1}{|x|^{p}}=\dfrac{1}{(p-2)(p-d)}\,\partial^{2}\,\dfrac{1}{|x|^{p-2}}~,$ (24)
$(ii)$ $\quad\mu\dfrac{d}{d\mu}\,\left({\cal{R}}\;\dfrac{1}{|x|^{d}}\right)=S_{d}\,\delta^{d}(x)~,$ (25)
it is direct to iterate with (i) and lower the exponent until one reaches $d$, and then (ii) produces the Dirac delta. The outcome of this iteration is
$\mu\dfrac{d}{d\mu}\langle T_{ab}(x)\,T_{cd}(0)\rangle=\dfrac{(d-1)\,(d/2-1)!}{(4\pi)^{d/2}\,(d+1)!}\,C_{T}\,\Delta^{T}_{ab,cd}\,(\partial^{2})^{d/2-2}\,\delta^{d}(x)~.$ (26)
On the RHS, in turn, we need the kernel corresponding to the second metric variation of the trace anomaly. It suffices to keep track of the quadratic term in the Weyl tensor. Furthermore, the Euler term does not contribute. One can directly expand the term quadratic in Weyl up to second order in the metric perturbation around flat space, or trade it for the Euler term and $Q$-curvature modulo total derivatives. The $Q$-curvature is pure Ricci and can be easily varied in 4D and 6D, where explicit expressions for it are at hand. Cubic and higher-order terms in the Weyl tensor can safely be discarded when expanding around flat space in the computation of the two-point function, since their contribution to the effective action is cubic or higher in the metric perturbation.
It may prove instructive to note that the connection between $C_{T}$ and $c$ can also be obtained within dimensional regularization by careful examination of the pole structure. On the $C_{T}$ side, we have that the kernel $1/|x|^{2d-4}$ has a pole, in a distributional sense  [^5] See e.g. the classic book by Gelfand and Shilov on generalized functions[52], given by
$\dfrac{1}{|x|^{2d-4}}\,\sim\,\dfrac{1}{\epsilon}\cdot\dfrac{(d-2)\,S_{d}}{2^{d-3}\,(d-3)!}\,(\partial^{2})^{d/2-2}\,\delta^{d}(x),$ (27)
that must be cancelled by the one-loop counterterm, which in turn, is given by the trace anomaly (see e.g. [26], eqns. 21-26 therein for 2D and 4D  [^6] We are grateful to M.J. Duff for information on this point. This same pole matching led to the discovery of the GJMS operators in the two-point function for scalar operators in [53]. See also  [54] for a thorough analysis of scalar operators.).
III.0.1 Standard derivation in 4D
In 4D, we need only expand Weyl’s conformal gravity $W^{2}$ on the RHS to quadratic order in the metric perturbation around flat space [1]. It is enough to consider the linear variation of the Weyl tensor that is given by
$W_{abcd}=2{\cal E}^{W}_{abcd,efgh}\,\partial_{f}\partial_{g}\,h_{eh}\,+O(h^{2}),$ (28)
where ${\cal E}^{W}_{abcd,efgh}$ denotes the projector onto the space of tensors with the same algebraic properties as the Weyl tensor ${\cal E}^{W}_{abcd,efgh}=\partial W_{abcd}/\partial W_{efgh}$. After some algebraic manipulations, the RHS can be cast in the form
$\mu\dfrac{d}{d\mu}\langle T_{ab}(x)\,T_{cd}(0)\rangle=\dfrac{C_{T}}{(4\pi)^{d/2}\,40}\,\Delta^{T}_{ab,cd}\,\delta^{4}(x)=\dfrac{4\,c}{(4\pi)^{2}}\,\Delta^{T}_{ab,cd}\,\delta^{4}(x)~,$ (29)
leading to the 4D relation
$C_{T}=160\,c.$ (30)
III.0.2 Alternative derivation in 4D
In 4D, we can take advantage of the fact that, modulo the topological Euler density, one can trade Weyl-squared conformal gravity by Lanczos conformal gravity, which is pure Ricci and essentially the critical 4D $Q$-curvature, modulo a trivial total derivative, by virtue of the Gauss-Bonnet formula. This path was undertaken by Sen and Sinha [28], who expanded the Ricci tensor and scalar on Cartesian coordinates
$R_{ab}=-\frac{1}{2}\left(\partial^{c}\rho\partial_{a}h_{bc}+\partial^{c}\partial_{b}h_{ac}-\partial^{2}h_{ab}-\partial_{a}\partial_{b}h\right)+O(h^{2})\text{ and }R=-\partial^{a}\partial^{b}h_{ab}+\partial^{2}h+O(h^{2}),$ (31)
in order to derive the second-order term in the metric variation around a flat space of
$W^{2}\sim 2(Ric^{2}-R^{2}/3)\sim-4\,Q_{4},$ (32)
and obtained the very same result as before.
III.0.3 Standard derivation in 6D
Here we need to expand to quadratic order in the metric fluctuation $h_{ab}$ the six-dimensional Weyl conformal gravity $I_{3}=W\nabla^{2}W+lot+ttd$ around flatspace, resulting in [29]
$\mu\dfrac{d}{d\mu}\langle T_{ab}(x)\,T_{cd}(0)\rangle=\dfrac{C_{T}}{(4\pi)^{3}\,504}\,C_{T}\,\Delta^{T}_{ab,cd}\partial^{2}\,\delta^{6}(x)=\dfrac{6\,c_{3}}{(4\pi)^{3}}\,\Delta^{T}_{ab,cd}\partial^{2}\,\delta^{6}(x)~,$ (33)
leading to the 6D relation
$C_{T}=3024\,c_{3}.$ (34)
III.0.4 Alternative derivation in 6D
The six-dimensional analog of Lanczos conformal gravity is given by the Lü, Pang, and Pope conformal gravity [30], written in terms of the Ricci tensor and its derivatives. This 6D conformal gravity is exactly the holographic Weyl anomaly of [9] and, after Graham and Zworski, the 6D critical $Q$-curvature. It is enough to restrict attention to the term quadratic in curvature, so that, up to the Euler density, cubic in curvature terms and trivial total derivatives, we have
$I_{3}\sim W\nabla^{2}W\sim 3\,Ric\nabla^{2}Ric-\dfrac{9}{10}R\nabla^{2}R\sim 6\,Q_{6}\sim\dfrac{3}{2}W^{(6)}_{2,1}$. (35)
After some algebraic manipulations while expanding the Ricci tensor and scalar, as expected, one obtains the very same relation as before  [^7] This 6D relation was announced by Sen and Sinha [28], but it contained a wrong numerical coefficient in the holographic computation of the central charge $c_{3}$ that spoiled the relation with $C_{T}$ in the journal version. The arXiv version 4, in turn, contains the correct result that agrees with the holographic expectation; see eqns. 3.22 and D.29 therein. We are grateful to Felipe Diaz for clarification of this issue..
III.0.5 Standard derivation in general even dimensions
Let us now present both derivations in a generic even dimension. We need to expand the term quadratic in the Weyl tensor in the trace anomaly. Within our conventions, this corresponds, up to trivial total derivatives and higher order terms in curvature, to the term
$(4\pi)^{d/2}\langle T\rangle=c\,\frac{d(d-2)}{8(d-3)}\,W\left(\nabla^{2}\right)^{d/2-2}W+lot+ttd\quad.$ (36)
Conformal symmetry fixes the structure of the RHS except for the numerical coefficient #\# in front of the $c$ charge
$\mu\dfrac{d}{d\mu}\langle T_{ab}(x)\,T_{cd}(0)\rangle=\dfrac{(d-1)\,(d/2-1)!}{(4\pi)^{d/2}\,(d+1)!}\,C_{T}\,\Delta^{T}_{ab,cd}\,(\partial^{2})^{d/2-2}\,\delta^{d}(x)=\dfrac{\#\,c}{(4\pi)^{d/2}}\,\Delta^{T}_{ab,cd}(\partial^{2})^{d/2-2}\,\delta^{d}(x)~.$ (37)
Let us take a shortcut to avoid gymnastics with the many indices. If we restrict ourselves to acting on the transverse-traceless part of the metric fluctuation $h$ and go to momentum space, we can employ a useful identity worked out by Erdmenger and Osborn, see [32], together with the fact that ${\cal E}^{W}_{abcd,efgh}$ is a projector. Back to position space, the quadratic expansion reads
$W\left(\nabla^{2}\right)^{d/2-2}W=\frac{d-3}{d-2}\,h^{tt}_{ad}\,(\partial^{2})^{d/2}\,h^{tt}_{ad}\,+\,O(h^{3}).$ (38)
Functional derivatives produce a factor of 2 and two more factors of 2 from the definition of the energy-momentum tensor, and the quartic operator $\Delta^{T}_{ab,cd}$ on transverse traceless tensors simply produces $(\partial^{2})^{2}$ times identity. So, the missing numerical factor is simply #=d\#=d. Comparing both sides of the equality, we arrive within the CFT to the universal relation
$\boxed{C_{T}\;=\;\frac{d}{d-1}\cdot\frac{(d+1)!}{\left(d/2-1\right)!}\cdot c}$ (39)
III.0.6 Alternative derivation: $Q$-curvature and GJMS-like operators
As for the alternative route, we can invoke the connection with the $Q$-curvature, a pure Ricci curvature invariant. The metric variation of the critical $Q$-curvature is the obstruction tensor, and its leading asymptotics is generally known  [^8] Our thanks to J.S. Case for the reference and further clarifications.
$Q_{d}$ $=$ $(-\nabla^{2})^{d/2-1}\dfrac{R}{2(d-1)}\,-\,\dfrac{d-2}{8(d-3)}\,W(-\nabla^{2})^{d/2-2}W\,+lot$ (40)
$=$ $(-\nabla^{2})^{d/2-1}\dfrac{R}{2(d-1)}\,-\,\left\{\dfrac{1}{2}\,Ric(-\nabla^{2})^{d/2-2}Ric\,-\,\dfrac{d}{8(d-1)}R(-\nabla^{2})^{d/2-2}R\right\}\,+lot.$ (41)
We can therefore trade the term quadratic in Weyl by the $Q$-curvature
$(4\pi)^{d/2}\langle T\rangle=-c\,(-1)^{d/2}\,d\,Q_{d}\,+lot+ttd\quad.$ (42)
As mentioned above, the metric variation of the $Q$-curvature is the Fefferman-Graham obstruction tensor (the higher-dimensional analog of the 4D Bach tensor), so we only need the linearization of the obstruction tensor around flat space.
But there is a remarkable observation following this line of reasoning: the second metric variation of the $Q$-curvature, which can be thought of as a particular conformal gravity Lagrangian admitting Einstein solutions, is just the kinetic term of the corresponding Weyl graviton when restricted to the transverse-traceless component of the metric fluctuation. This is nothing but the GJMS-like operator on symmetric transverse-traceless two-tensors that happens to factorize into shifted Lichnerowicz Laplacians on Einstein manifolds, as shown some time ago by Matsumoto [34]. In particular, in flat space, it reduces to a power of the Laplacian. It is just a matter of keeping track of the overall coefficient to be able to obtain the numerical factor that fixes the RHS of eqn.(37)  [^9] We are grateful to A.A.Tseytlin for discussion of this point.
$Q_{d}=-\frac{1}{8}\,h^{tt}_{ad}\,(-\partial^{2})^{d/2}\,h^{tt}_{ad}\,+\,O(h^{3}).$ (43)
That completes the alternative derivation à la Lanczos, i.e., trading Weyl (or Riemann) tensors by pure Ricci terms in the conformal gravity Lagrangian.
IV Examples
There is a vast literature on four and six-dimensional CFTs [36, 37] and references therein. We will focus on the less-explored instances $D\geq 8$.
For conformal powers of the Laplacian, GJMS operators, there was a conjectured $C_{T}$ value for up to the critical one (i.e. $(\partial^{2})^{d/2}$ in flat space) due to Osborn and Stergiou [38]. In their original paper, the authors verified the results in two, four, and six dimensions. The eight-dimensional case was confirmed shortly thereafter in [39], providing some insight into a general proof, based on Conformal Partial Wave expansions of four-point functions for free fields that they summed over conformal primaries.
Field $C_{T}$
$\partial^{2}$ conformal scalar d/(d1)d/(d-1)
$\partial^{4}$ conformal scalar 2d(d+4)/(d1)(d2)-2d(d+4)/(d-1)(d-2)
$\partial^{6}$ conformal scalar 3d(d+4)(d+6)/(d1)(d2)(d4)3d(d+4)(d+6)/(d-1)(d-2)(d-4)
$\partial^{8}$ conformal scalar 4d(d+4)(d+6)(d+8)/(d1)(d2)(d4)(d6)-4d(d+4)(d+6)(d+8)/(d-1)(d-2)(d-4)(d-6)
8D GJMS
Taking advantage of the factorization properties of the GJMS operators on Einstein manifolds, we have been able to compute the accumulated relevant heat coefficient by using the explicit result of Avramidi[40] for the $b_{8}$ diagonal heat coefficient in 8D. The boundary anomaly was also reproduced holographically from the heat coefficients in 9D of the dual bulk massive scalar [41], extending the program initiated in [42] for 5D/4D and 7D/6D. The $c$ central charge for the 8D $P_{2k}$ operators is given by the following polynomial in $k$
$9!\cdot c=-\frac{k^{9}}{24}+\frac{3k^{7}}{4}-\frac{91k^{5}}{40}-\frac{41}{6}k^{3}+\frac{72k}{5}~.$ (44)
Comparison with the expectation of the Osborn-Stergiou predictions can only be made for subcritical and critical $P_{2k}$, i.e., for $k=1,2,3$ and $4$ in 8D. In the following, we tabulate results that are in complete agreement with the previously derived universal relation between $C_{T}$ and $c$.
$k$ Field $C_{T}$ $9!\cdot c$
11 $\partial^{2}$ conformal scalar $8/7$ 66
22 $\partial^{4}$ conformal scalar 32/7-32/7 24-24
33 $\partial^{6}$ conformal scalar 2424 126126
44 $\partial^{8}$ conformal scalar 256-256 1344-1344
Holographic higher-order gravities
We can also add to the list of examples, of course, the holographic data for bulk Einstein gravity that we used in the first place to obtain the numerical relation
$C^{E}_{T}=4\pi^{d/2}\cdot\frac{\Gamma(d+2)}{(d-1)\,\Gamma^{3}(d/2)}\cdot\frac{L^{d-1}}{l_{\text{P}}^{d-1}}\qquad vs.\qquad c^{E}=\dfrac{4\pi^{d/2}}{d\,\Gamma^{2}(d/2)}\cdot\dfrac{L^{d-1}}{l_{\text{P}}^{d-1}}~.$ (45)
Actually, we can do better and consider adding to the EH Lagrangian any (local) higher curvature term, i.e. contractions of the Riemann tensor and including its covariant derivatives as well
[…]
here $L$ is a length scale introduced so as to make the coupling $\alpha_{n}$ dimensionless. The (Euclidean) AdS vacuum solution of the original EH action has radius $L$, but in general, we will have an effective radius for the putative AdS vacuum solution. At linear order in the couplings $\alpha_{n}$ it will exist, and in general one needs to find a positive (generically nondegenerate) root of an algebraic equation for the ratio between the parameter $L$ and the effective AdS radius $\tilde{L}=L/\sqrt{f_{\infty}}$.
Let us follow a shortcut route [43, 44, 45] to get this algebraic equation . This is equivalent to extremizing the gravitational action with respect to the AdS radius as in [7]. We exploit the scale invariance of the action to write
[…]
and then take
$\dfrac{d}{d\mu}S_{G}[\mu^{2}\,\hat{g}]{|}_{\mu=1}=0.$ (48)
It is immediate to realize that on the AdS vacuum, by dimensional arguments, the Lagrangian must depend on the ratio $f_{\infty}/\mu^{2}$
$S_{\text{G}}[\mu^{2}\hat{g}_{AdS}]=\mu^{d+1}\int_{\mathcal{M}}d^{d+1}x\sqrt{\hat{g}_{AdS}}\;{\cal L}\left(\frac{f_{\infty}}{\mu^{2}}\right),$ (49)
so that the scale invariance condition translates into
$(d+1)\,{\cal L}(f_{\infty})\,-\,2\,f_{\infty}\,{\cal L}^{\prime}(f_{\infty})=0.$ (50)
It is convenient to rewrite this algebraic equation in terms of the auxiliary function
$h(f_{\infty})=\frac{-2l_{\text{P}}^{d-1}\,L^{2}}{d(d-1)}\left[{\cal L}(f_{\infty})\,-\,\frac{2\,f_{\infty}}{d+1}\,{\cal L}^{\prime}(f_{\infty})\right],$ (51)
that on-shell satisfies $h(f_{\infty})=0$ and has the simple expansion, after suitable rescaling the invariants,
$h(f_{\infty})=1-f_{\infty}\,+\,\sum_{n\geq 2}\alpha_{n}f_{\infty}^{n}.$ (52)
We can now apply our simple holographic recipe [45, 41] to determine the holographic central charges $a$ and $c$ by evaluating the action on the particular Poincaré-Einstein bulk metric with an Einstein metric at its conformal infinity. From the pure-Ricci part, we get the action at $AdS$ precisely, and from the deviations, involving the Weyl tensor, it is enough to keep the Weyl-squared term
$S_{\text{G}}[\hat{g}_{{}_{PE/E}}]={\cal L}(f_{\infty})\cdot\int_{\mathcal{M}}d^{d+1}x\sqrt{\hat{g}_{{}_{PE/E}}}\quad+\quad\int_{\mathcal{M}}d^{d+1}x\sqrt{\hat{g}_{{}_{PE/E}}}\cdot\gamma\,\hat{W}^{2}\quad+\quad\cdots.$ (53)
From the first term we obtain the volume anomaly, proportional to the critical $Q$-curvature $Q_{d}$, and for the quadratic $\hat{W}^{2}$ term we complete with derivatives to obtain the particular conformal invariant on Einstein manifolds $\hat{W}_{2,1}^{(d+1)}$, quadratic in the bulk Weyl tensor, as obtained by Case et al.[13], that induces the corresponding Weyl invariant $W^{(d)}_{2,1}$ on the Einstein boundary. Keeping track of the numerical coefficients, we obtain for the holographic Weyl anomaly
$(4\pi)^{d/2}\langle T\rangle$ $=$ $\,-(-1)^{d/2}\cdot a\cdot\,d\,Q_{d}+\,(c-a)\dfrac{d}{4}\,W^{(d)}_{2,1}\,+lot+ttd\quad.$ (54)
$=$ $-(-1)^{d/2}\cdot\dfrac{\pi^{d/2}\,\tilde{L}^{d+1}}{[(d/2)!]^{2}}\cdot{\cal L}(f_{\infty})\cdot\,d\,Q_{d}\,-\,\dfrac{(2\pi)^{d/2}\,\tilde{L}^{d-3}}{(d/2)!\,(d-4)!!}\cdot 8\gamma\cdot\dfrac{d}{4}\,W^{(d)}_{2,1}\,+lot+ttd\quad.$ (55)
In all, the holographic central charges $a$ and $c$ are given by
$a$ $\,=\,$ $\dfrac{\pi^{d/2}\,\tilde{L}^{d+1}}{[(d/2)!]^{2}}\cdot{\cal L}(f_{\infty}),$ (56)
$c-a$ $\,=\,$ $-\,\dfrac{8d(d-2)\,\pi^{d/2}\,\tilde{L}^{d-3}}{[(d/2)!]^{2}}\cdot\gamma.$ (57)
This result for the central charge $a$ is well known, first obtained by Imbimbo et al. [46]. As for the $c$ charge, in the particular case of Einstein‑like gravities—though this still comprises a large family of gravitational theories (cf. [47])—the structure of the coefficient $\gamma$ admits further refinement
$L^{4}\,{\cal L}^{\prime\prime}(f_{\infty})=2(d+1)d(d-1)(d-2)\,\gamma,$ (58)
and, after basic algebraic steps, we derive an expression for $c$ that tweaks the canonical Einstein–Hilbert value in a notably simple way
$c\,=\,-h^{\prime}(f_{\infty})\cdot\dfrac{d\pi^{d/2}\cdot\tilde{L}^{d-1}}{[(d/2)!]^{2}\cdot l_{\text{P}}^{d-1}}\,=\,-h^{\prime}(f_{\infty})\cdot c^{E}|_{\tilde{L}}$. (59)
Remarkably, using our universal relation, we can confirm the corresponding value for $C_{T}$ reported in [47]
$C_{T}\,=\,-h^{\prime}(f_{\infty})\cdot\frac{4\pi^{d/2}(d+1)!\cdot\tilde{L}^{d-1}}{(d-1)\,[(d/2-1)!]^{3}\cdot l_{\text{P}}^{d-1}}\,=\,-h^{\prime}(f_{\infty})\cdot C^{E}_{T}|_{\tilde{L}}$. (60)
Let us finish this discussion by recalling the observation of Li, Lü and Mai [48] on $a$ and $c$. Provided $a$ and $c$ are written in terms of the effective AdS radius $\tilde{L}$ on-shell, i.e., with L=L(L~)L=L($\tilde${L}) and $f_{\infty}$ being a mere number determined by the theory’s couplings, then it is satisfied
$c\,=\,\dfrac{1}{d-1}\,\tilde{L}\,\dfrac{\partial}{\partial\,\tilde{L}}\,a.$ (61)
This can be readily verified by taking the (log-)derivative of the holographic central charge $a$ (eqn.56)
$\tilde{L}\,\dfrac{\partial}{\partial\,\tilde{L}}\,a\,$ $=$ $\,\dfrac{2\pi^{d/2}\,L^{d+1}}{(d+1)\,[(d/2)!]^{2}}\,\left[-2\,f_{\infty}\dfrac{\partial}{\partial\,f_{\infty}}\,\right]\left\{f_{\infty}^{-(d-1)/2}\cdot{\cal L}^{\prime}(f_{\infty})\right\}$ (62)
$=$ $\,-h^{\prime}(f_{\infty})\cdot\dfrac{d(d-1)\pi^{d/2}\cdot\tilde{L}^{d-1}}{[(d/2)!]^{2}\cdot l_{P}^{d-1}}\,=\,-h^{\prime}(f_{\infty})\cdot(d-1)\cdot c^{E}|_{\tilde{L}}\,=\,(d-1)\cdot c,$ (63)
where we have used the on-shell condition to trade ${\cal L}(f_{\infty})$ by $\frac{2\,f_{\infty}}{d+1}\,{\cal L}^{\prime}(f_{\infty})$ and the fact that the latter only depends on $\tilde{L}$, as opposed to ${\cal L}(f_{\infty})$ which also depends on $L$. That makes it possible to act with derivatives with respect to $f_{\infty}$. Using the present universal relation, we end up with a prescription to compute $C_{T}$ directly from the type-A central charge $a$ in Einstein-like gravities. A comparison with [47] (Eq.(18) and footnote [90] therein) is also in order.
V Conclusion
In summary, we unveiled the universal relationship between the CFT coefficient $C_{T}$ and the central charge $c$ in arbitrary even dimensions. This highlights the nontrivial connection between conformal field theories in flat space and the trace anomaly arising from gravitational coupling. We have presented three different derivations, a holographic one and two genuine CFT ones. Perhaps the most concise –and most striking– formulation of the universal relation is in terms of the coefficient $C_{TT}$ of the momentum-space TT-correlator (as flat as it gets) and the critical Q-curvature $Q_{d}$ (as curved as it is)
$\langle T\rangle\,=\,2\,C_{TT}\cdot Q_{d}+lot+ttd.$ (64)
This effectively extends to higher-even dimensions previously known results [1, 32, 28, 38] and, at the same time, contains the Weyl-completion of the term quadratic in the Weyl tensor (cf. [23], Eq. (3.17)).
Notably, several other contexts have established a similar connection to $C_{T}$. These include
  • Rényi entropy across a spherical entangling surface [49]
  • Free energy on conically deformed spheres [50]
  • Free energy on squashed spheres [47]
  • Pseudoentropy for (slightly) deformed spheres in dS/CFT [51]
A direct mapping between the above quantities and the Weyl anomaly coefficients appears compelling.
In the interest of universality, it would also be instructive to consider the energy flux parameters $t_{2}$ and $t_{4}$. These parameters govern the three‑point function of the energy–momentum tensor and are expected to be related to the Weyl‑cubic contributions to the trace anomaly in six or more dimensions.
VI Acknowlegments
We are grateful to G.Anastasiou, J.S.Case, F.Diaz, M.J.Duff, P.McFadden, A.A.Tseytlin and O.Zanusso for useful discussions, clarifications, and for pointing out relevant references. DED would also like to thank the faculty and staff of the Science Department at Valencia College for their warm welcome and support. This work was partially funded through FONDECYT-Chile 1220335. RA wishes to thank professors J. Gomis (U. Barcelona), E. Bergshoeff (U. Groningen), M. Romo and H. Adami (SIMIS, Shanghai), and E. Joung (U. Kyung Hee, Seoul) for their kind hospitality during the completion of this work.
Note added: After completion of the first version of this work, we were informed by P. McFadden that in [23], the authors carried out the equivalent to our standard CFT derivation in momentum space and were able to relate their $C_{TT}$ to the coefficient of the $W(\nabla^{2})^{d/2-2}W$ term in the trace anomaly. Their result turns out to be correct in $4D$, but the appropriate extension to higher even dimensions requires an additional dimensional factor (d2)/2(d3)(d-2)/2(d-3) that escaped notice. However, we find their relation very instructive and, consequently, we have cast our own result in terms of $C_{TT}$, absorbing the dimensional factor into the critical Q‑curvature $Q_{d}$, thereby giving a new turn to the universal relation.